OpenStudy (anonymous):

What number is 16% of y use a proportion

OpenStudy (asnaseer):

remember that percentage means "out of a hundred". so 16% is 16/100 which can be simplified

OpenStudy (anonymous):

4y/25

hero (hero):

\[\frac{x}{y} = \frac{16}{100}\]

hero (hero):

said use PROPORTION

hero (hero):

100x = 16y x = 16y/100 x = 4y/25